OpenStudy (anonymous):

A pet shop had 8 identical fish tanks in the shape of rectangular prisms. The width and height of each tank was (x - 7) inches and the length was 8 inches more than its width. The expression below shows the maximum amount of water which all the tanks together can hold. 8(x + 1)(x - 7)2

OpenStudy (anonymous):

(x + 1)(x - 7)^2 is the volume of EACH tank.

OpenStudy (anonymous):

The volume of all the tanks 8 identical tanks Volume of each tank; L*W*H = (x-7+8)(x-7)(x-7) = (x+1)(x-7)^2 so, for 8 tanks: 8(x+1)(x-7)^2